Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing difficulty for some time. The company’s contribution format income statement for the most recent month is given below:

Total (Tk.)

Per Unit

Percent of Sales

Sales (19,500 units)

5,85,000

30

100%

Variable Expense

   4,09,500

21

? %

Contribution Margin

   1,75,500

9

? %

Fixed Expenses

   1,80,000

Net Operation Income

(4,500)

Required:

  1. Compute the company’s CM ratio and its break-even point in both units and dollars.
  2. The president believes that a $16,000 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $80,000 increase in monthly sales.
  3. If the president is right, what will be the effect on the company’s monthly net operating income or loss? (Use the incremental approach in preparing your answer.)
  4. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $60,000 in the monthly advertising budget, will cause unit sales to double. What will the new contribution format income statement look like if these changes are adopted?
  5. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would help sales. The new package would increase packaging costs by 75 cents per unit. Assuming no other changes, how many units would have to be sold each month to earn a profit of $9,750?
  6. Refer to the original data. By automating certain operations, the company could reduce variable costs by $3 per unit. However, fixed costs would increase by $72,000 each month.
  1. Compute the new CM ratio and the new break-even point in both units and dollars.
  2. Assume that the company expects to sell 26,000 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.)
  3. Would you recommend that the company automate its operations? Explain.
Answer #1

Adjustment 1.

Total

Per unit

Percent of sales

Sales (12800 units)

585,000

$ 30 per unit

100%

Variable expenses

409,500

21

70

Contribution margin

175,500

9

30%

Contribution margin ratio is 30%

Break-even point shall be:

Unit sales to breakeven = Fixed expenses/unit contribution margin = 180,000/9 = 20,000 units

Dollar sales to breakeven = Fixed expenses/CM ratio = 180,000/0.3 = $ 600,000

Adjustment 2.

$80,000 increased sales X Contribution ratio 30%

24,000

Less: increased advertising cost

16000

Increase in net monthly operating income

8000

Currently the company is making loss of 4,500 which shall be increased by 8,000 when the sales are increased. The profit would turn into 8000-4500 = $3,500

Adjustment 3.

New selling price = 30 -10% of 30 = $ 27

Sales (19500 *2 =39000) = 39000*27

1053,000

Variable expenses (39000*21)

819,000

Contribution margin

234,000

Fixed expenses (180000+60000)

240,000

Operating gain

(6000)

Adjustment 4.

New contribution margin = 9- 0.75 = 8.25

Unit sales to attain target profit = (Target profit +Fixed expenses)/ CM per unit

                                                                = (9750+180,000)/8.25

                                                                = 23000 units

Adjustment 5.

Reducing variable expenses by $ 3 and increasing fixed cost by 72,000

Total

Per unit

Percent of sales

Sales (19500 units)

585000

$ 30 per unit

100%

Variable expenses

351000

21-3 = 18

60%

Contribution margin

234000

12

40%

a.

Contribution margin ratio is 40%

Break-even point shall be:

Unit sales to breakeven = Fixed expenses/unit contribution margin = (180000+72,000)/12 = 21000 units

Dollar sales to breakeven = Fixed expenses/CM ratio = (180000+72000)/0.4 = $ 630,000

b.

Contribution margin statement

For 26,000 units

Non-automated

Automated

Total

Per unit

%

Total

Per unit

%

Sales

780,000

30

100

780,000

30

100

Variable expenses

546,000

21

50

468,000

18

60

Contribution margin

234,000

9

50

312,000

12

40

Fixed expenses

180,000

252,000

Net operating income

54,000

60,000

c. Automation should be recommended as there is more net operating income than non-automated